OpenStudy (burberry22):

Problem Page A line passes through the point , −4−2 and has a slope of −3 . Write an equation in point-slope form for this line.

OpenStudy (akshvas):

y-y1 = m*(x-x1) where m is the slope and x1, y1 is the coordinate on the line.

OpenStudy (akshvas):

Plug in the values to get the answer

OpenStudy (burberry22):

oh ok

OpenStudy (akshvas):

Do you know how does one derive to that form of equation? It's called the point slope form.

OpenStudy (akshvas):

The derivation of that form comes from the basic definition of slope.

Directrix (directrix):

Point-slope form is written like this: (y - y1) = m(x - x1) In point-slope form, y1 is the y value of the known point on the line, m is the slope and x1 is the x value of the known point. In your problem, (x1,y1) = (−4−2) m= −3 Fill that into the form for the line: (y - y1) = m(x - x1)

OpenStudy (burberry22):

i have a huge headache can you give m the answer bro?

OpenStudy (akshvas):

3x + y + 14 = 0